# The Surprising Soupability of Documents in State Space Models

### v1

@@ 3.1.2 Evaluation Metrics
For multiple-choice QA tasks, including RACE and QuALITY, we use multiple-choice accuracy as the evaluation metric. In this setting, the model outputs a distribution over the four answer options (A, B, C, D), and we compute the log-probability of each choice. The predicted answer corresponds to the option with the highest score, and accuracy is measured as the proportion of correctly selected answers.
3.1.3 Baselines and Reference Setting
We compare souping against several baselines that represent different levels of supervision and context integration. First, we include a pretrained baseline: an off-the-shelf Mamba2 model used without any task-specific finetuning. This setup serves as a lower bound, as the model has not been adapted to the QA tasks.
Next, we consider a QA-only finetuning baseline. Here, the model is trained only on $(q,a)$ pairs without access to any document context during training. Although this configuration does not support long-context integration, it provides a useful reference for how much performance is attributable to the context.
Finally, we evaluate a concat-based finetuning baseline, where the model is trained end-to-end on full input sequences—i.e., $(d_{1},\dots,d_{k},q,a)$—concatenated into a single flat input. This setting provides strong supervision by exposing the model to all relevant documents in a joint context window and serves as our primary reference point when assessing the performance of souping-based alternatives.
Our aim is to demonstrate that, with appropriate finetuning, souping-based models can outperform context-free baselines and approach the performance of the concat-based reference model.
@@ 3.1.3 Baselines and Reference Setting
Soup w/ Average 2 63.62 59.43 56.49 52.30 47.46 47.27 45.16 43.10
4 66.02 63.54 60.58 57.31 51.15 50.05 48.08 45.69
8 62.29 59.28 55.74 53.77 46.31 45.69 44.20 43.29
3.2 Training Strategies for Soupability
We evaluate three training regimes that differ in how the model is exposed to and trained on soupable hidden states.
In the out-of-the-box setting, we merge hidden states from a pretrained Mamba2 encoder and pass them to the decoder without any finetuning. As expected, performance is poor, as the model has not learned to interpret pooled representations.
In the decoder-only finetuning setup, the encoder is frozen and only the decoder is trained to answer questions from fixed, aggregated states. This tests whether the decoder alone can learn to interpret pooled inputs. Performance improves substantially over the baseline but remains constrained by the unadapted encoder outputs.
Full encoder-decoder finetuning trains both components jointly, allowing the encoder to produce more mergeable states and the decoder to interpret them effectively. As shown in Table 1, this setup consistently achieves the strongest results, confirming that SSMs can be trained end-to-end to support modular reasoning through state pooling.
3.3 Aggregation Strategy Comparison
We evaluate several strategies for aggregating per-layer hidden states across documents, including elementwise average, sum, and maximum, along with optional unit normalization applied before or after aggregation. Results are reported in Table 1.
When evaluating a model trained on 5 documents and tested across varying soup sizes, we observe that overall performance is relatively stable across aggregation methods, and no-norm averaging consistently yields strong performance. This robustness suggests that averaging provides a natural balance: it aggregates across documents while keeping the scale of the hidden states stable.
By contrast, summation without normalization tends to degrade more quickly as more documents are merged. We hypothesize that this is due to the unbounded growth in activation magnitude. Applying normalization after summation improves performance at larger document counts. However, it still underperforms averaging, which implicitly maintains consistent magnitude while preserving directional information. Interestingly, normalization before aggregation performs well with averaging but poorly with summation, likely because it destroys relative magnitude differences across documents that are helpful when using sum.
Taken together, these observations suggest that averaging without normalization offers the best trade-off between simplicity, stability, and generalization. We adopt this as the default aggregation method in all subsequent experiments.
Method Train Test Segments
Segments 1 2 5 10 20
Concat 1 54.81 / 71.90
Soup w/ Average 2 58.38 / 74.05 34.89 / 49.04 13.19 / 23.40 12.36 / 22.45
5 58.38 / 73.89 57.28 / 72.09 35.71 / 50.80 14.97 / 26.25
10 21.02 / 31.41 13.87 / 22.36 52.75 / 68.68 43.82 / 58.99
20 28.71 / 41.44 28.85 / 42.93 45.60 / 61.78 44.37 / 60.85
3.4 Generalization Across Soup Sizes
We study how the number of documents or segments being merged at inference time affects model performance, particularly when this number differs from the training configuration. This analysis addresses the extent to which soupability generalizes across varying soup sizes.
Tables 2, 3, 4, and 5 show that performance typically peaks when the number of merged inputs at inference matches the training-time setting. Increasing the soup size beyond the training configuration leads to a gradual performance drop, likely due to information dilution or representational interference. Reducing the number of inputs typically causes a smaller degradation, suggesting some robustness to under-composition. Notably, models trained with larger soup sizes exhibit better generalization to even larger settings. For example, a model trained on 8 segments maintains strong performance when evaluated on 16 segments, whereas a model trained on only 2 segments generalizes poorly in this setting. This trend is consistent across both QA and retrieval tasks.
These results suggest that exposure to more fragmented input during training helps the model learn representations that are more tolerant to variation in soup size. When paired with appropriate finetuning, this property enables robust reasoning over multi-document contexts of varying width.
3.5 Effect of Corpus Structure on Soupability
We evaluate how different input structures—specifically long single-document narratives and multi-document corpora—affect the effectiveness of state souping.
Long Documents.
To assess whether souping is viable for long, contiguous inputs, we experiment with the RACE and QuALITY datasets. RACE consists of relatively short reading passages, while QuALITY includes significantly longer and more complex narratives. Both are evaluated in a multiple-choice QA setting.
Interestingly, models trained with souping on RACE generalize well to QuALITY (Table 4), despite the large increase in input length. This is likely due to the structure of SSMs: each document’s hidden state is a fixed-size vector, independent of the input length. As a result, the model can continue to aggregate representations effectively, even as document or segment length increases. In contrast, models finetuned using standard concatenation on RACE generalize poorly to QuALITY, where the input sequence length grows beyond what the model was exposed to during training. This suggests that souping offers a more length-invariant mechanism for representation and generalization.
Multi-hop Reasoning.
We also examine whether souping supports compositional reasoning over multiple documents, as required by HotpotQA. As shown in Table 2, models trained with souping can perform multi-hop reasoning across independently encoded documents, achieving performance on par with concat-based models. This indicates that the merged hidden states preserve cross-document relationships without requiring joint encoding.
Together, these results highlight that souping is robust to input granularity—from short segments to long narratives—and supports both single-hop and multi-hop QA. This flexibility is critical for applications that must operate over heterogeneous corpora without the cost of re-encoding or sequence length tuning.
@@ 4 Related Work
Transformers [21] remain the dominant architecture for modeling long-range dependencies, but their $\mathcal{O}(L^{2})$ attention cost and $\mathcal{O}(L)$ memory growth in the KV cache hinder scalability. A rich body of work explores architectural modifications to overcome this, including sparse attention [4, 1, 29], linearized attention [12], and chunked processing with memory compression [5, 26, 27]. Retrieval-augmented generation methods such as RAG [15] provide complementary approaches by offloading long-context memory to external sources.
Structured State Space Models (SSMs) offer an alternative path with fundamentally different scaling characteristics. Models like S4 [10, 9] introduced linear-time computation via structured recurrences. Mamba [8] builds on this with input-dependent gating, while Mamba2 [6] unifies SSMs and attention through structured state space duality, yielding substantial speedups on long sequences with strong downstream performance. Recent work further demonstrates the utility of Mamba-based models for dense and retrieval-augmented tasks, including dense passage ranking [31] and long-document retrieval in RAG pipelines [3].
Model and Task Souping.
Model souping refers to merging parameters across finetuned checkpoints to improve robustness and generalization without retraining [25, 20]. Recent work has explored souping internal representations rather than weights. State Soup [17] linearly combines task-specific hidden states for skill transfer. In contrast, our method focuses on merging hidden states from disjoint document chunks—enabling compositional reasoning across distributed corpora through simple aggregation strategies.
@@
Efficient ingestion of long contexts has been addressed through sparse or structured attention mechanisms [4, 29], memory compression across chunks [5, 26], and retrieval-based pipelines [15]. For SSMs, constant-memory recurrence and linear compute make them naturally suited to chunk-wise streaming. However, to our knowledge, no prior work has explicitly studied post-hoc merging of hidden states across chunks for joint reasoning.
@@
Recurrent models have long supported state-passing across time steps, and early works explored implicit mixing of information through recurrent transitions [7]. Our approach differs in that it investigates explicit aggregation of intermediate hidden states produced in parallel, rather than temporal chaining of recurrent updates.
@@
Recent work has explored adaptive merging of key/value caches to compress context without retraining [24], and memory-efficient cache eviction policies [32]. While related in spirit, these approaches focus on attention-based caches, whereas our work targets SSM hidden states.
@@
In the vision domain, token merging has been studied as a means of compressing long sequences. TCFormer [30] clusters tokens across space for progressive downsampling. Other work merges tokens dynamically based on content [2, 13], or selects value/key pairs to retain or evict during inference [23]. These ideas echo our motivation to aggregate compact, composable representations—though in our case applied to document states within state space models.
@@ 5 Conclusion
We present document souping, a method for merging hidden states across independently encoded documents in structured state space models. By aggregating layer-wise representations through simple commutative operations, souping enables modular, parallel ingestion of long-context inputs while supporting accurate downstream inference. Our experiments demonstrate that Mamba2 models trained with souping are capable of multi-hop reasoning, sparse retrieval, and generalization across document and segment scales—often matching or outperforming traditional concat-based finetuning.
While the results are promising, our study has several limitations. First, we were unable to evaluate soupability at very large segment counts (e.g. 128) due to our compute and memory constraints; investigating the limits of scalability in high-fragmentation regimes remains an open question. Second, we did not benchmark training or inference efficiency directly. Such analysis would require large-scale deployment infrastructure to meaningfully measure throughput and memory usage under realistic workloads, which we leave to future work.
Overall, our findings highlight state souping as a lightweight and effective strategy for long-context reasoning in SSMs, and we hope it provides a foundation for broader adoption and further exploration.
@@ A.1 Data Formatting
Algorithm 1 Input Formatting for Concat- vs. Souping-data
1:Documents $\{d_{1},\ldots,d_{k}\}$, question $q$, answer $a$, flag $\mathrm{soup}\in\{\mathrm{true},\mathrm{false}\}$
2:Formatted input $x$
3:$x\leftarrow$ empty string
4:for $i=1$ to $k$ do
5:     if $\mathrm{soup}$ then
6:         $x\leftarrow x\mathbin{\oplus}d_{i}\mathbin{\oplus}\langle\mathrm{DOC\_SEP}\rangle$
7:     else
8:         $x\leftarrow x\mathbin{\oplus}d_{i}\mathbin{\oplus}\texttt{" "}$
9:     end if
10:end for
11:$x\leftarrow x\mathbin{\oplus}q\mathbin{\oplus}\texttt{" "}\mathbin{\oplus}a\mathbin{\oplus}\langle\mathrm{eos}\rangle$
@@ A.3 Pseudo-code
1:Document set $\{d_{1},\dots,d_{k}\}$, query $q$
2:Predicted answer $\widehat{y}$
3:Initialize empty list of layerwise states $\{H^{(l)}\}_{l=1}^{L}$
4:for each document $d_{i}$ do
5:     Compute hidden states: $\{h_{i}^{(1)},\dots,h_{i}^{(L)}\}\leftarrow\mathrm{SSM}_{\theta}(d_{i})$
6:     for $l=1$ to $L$ do
7:         Optionally normalize: $\tilde{h}_{i}^{(l)}\leftarrow h_{i}^{(l)}/\|h_{i}^{(l)}\|$
8:         Append $\tilde{h}_{i}^{(l)}$ to $H^{(l)}$
9:     end for
10:end for
11:for $l=1$ to $L$ do
12:     Pool document states: $h_{\text{soup}}^{(l)}\leftarrow\mathrm{pool}(H^{(l)})$
13:     Optionally normalize: $h_{soup}^{(l)}\leftarrow h_{soup}^{(l)}/\|h_{soup}^{(l)}\|$
14:end for
15:Predict: $\widehat{y}\leftarrow\mathrm{SSM}_{\theta}\left(q\mid\{h_{\text{soup}}^{(l)}\}_{l=1}^{L}\right)$
A.4 Decoding Hyperparameters
[…]
Soup achieved best validation performance with temperature $=0.3$, top-$p=0.5$, and top-$k=20$, while Concat performed optimally with temperature $=0.5$, top-$p=0.95$, and top-$k=30$. All subsequent experiments adopted these respective configurations for each method.
Table 6: HotpotQA results for Mamba-2 2.7B with $2$ gold documents and $n-2$ distractors. Cells show EM / F1. Underline marks models tested on the same docs they are trained on.
Model Evaluation: $2$ gold + $(n{-}2)$ distractors

### v2

@@ 3.3 Baselines and Reference Setting
4 Results
We present experimental results demonstrating that with proper finetuning, SSMs can effectively merge independently encoded documents for complex reasoning tasks. We establish this across four key findings: we first show that finetuning is essential to unlock this “soupability”, then identify simple averaging as the most robust aggregation method, and finally demonstrate the approach’s scalability, generalization, and significant inference efficiency.
4.1 SSM Finetuning Unlocks Document Soupability
Finetuning is crucial for unlocking soupability in pretrained SSMs. We systematically evaluate three training strategies on HotpotQA, with results presented in Table 1. Our findings demonstrate a clear progression in performance as more components of the model are adapted to the task of interpreting souped states. First, we test an out-of-the-box approach by merging hidden states from a pretrained Mamba2 model without any finetuning. As shown in the top rows of Table 1, this yields poor performance (e.g., 2.6 EM on 5 documents), confirming that the model cannot interpret pooled representations naively. Second, we evaluate a decoder-only finetuning strategy. This results in a substantial performance gain (38.8 EM on 5 documents), demonstrating that the decoder can learn to interpret the fixed representations produced by an unadapted encoder. Third, full encoder-decoder finetuning trains both components jointly. This approach consistently achieves the strongest results (47.8 EM on 5 documents), outperforming the decoder-only method. This indicates that optimal performance is achieved when the encoder is also trained to produce more effectively mergeable states. These findings confirm that SSMs can be trained end-to-end to support powerful, modular reasoning through state pooling.
4.2 Averaging Document Representations Improves Soupability
We evaluate several strategies for aggregating per-layer hidden states, including elementwise average, sum, and maximum, with optional unit normalization. The results, detailed in Table 1, demonstrate that no-norm averaging is consistently the most effective and stable approach. This method achieves a top score of 47.8 EM and 61.3 F1 on 5 documents. We hypothesize that this stability stems from the ability to aggregate information across documents while inherently keeping the magnitude of the resulting state vector from growing with the number of inputs. In contrast, other methods are less robust, particularly as the number of documents grows. For example, summation without normalization degrades sharply from 44.2 EM on 5 documents to just 25.0 EM on 10 documents. We attribute this to the unbounded growth in activation magnitude. While applying normalization after summation mitigates this collapse, recovering the score to 34.7 EM, it still underperforms the simple averaging result of 38.7 EM on 10 documents. Interestingly, normalization prior to summation performs poorly, likely because it destroys relative magnitude differences across documents that are helpful when using sum. Given these findings, we conclude that averaging without normalization offers the best trade-off between simplicity, stability, and generalization. We therefore adopt it as our default pooling strategy for all subsequent experiments.
Method Train Test Segments
Segments 1 2 5 10 20
Concat 1 54.81 / 71.90
[…]
10 21.02 / 31.41 13.87 / 22.36 52.75 / 68.68 43.82 / 58.99
20 28.71 / 41.44 28.85 / 42.93 45.60 / 61.78 44.37 / 60.85
4.3 Document Souping is Scalable and Generalizes to Wide Contexts
Our analysis reveals two key principles that govern how document souping generalizes across different context sizes, demonstrated consistently across a range of diverse tasks (Tables 2, 3, 5, and 4). First, performance is usually maximized when the inference-time soup size aligns with the training configuration. This trend is clearly visible in our long-document QA results (Table 5). For instance, on the QuALITY benchmark, a model trained on 4 document segments achieves 50.05% accuracy when tested on 4 segments, and as the number of test segments increases to 8 and 16, it gradually declines to 48.08% and 45.69%, respectively.
Second, models finetuned on a greater number of segments exhibit more robust generalization. This is demonstrated in our Needle-in-a-Haystack experiments (Table 3). A model trained on only 2 segments with 4k sequence length fails completely when tested on 16 segments with the same sequence length (2.3 EM), while a model trained on 8 segments maintains a strong EM score of 71.45% when tested on 16. This shows that exposing the model to wider contexts during finetuning is key for robust generalization.
To stress-test the limits of this generalization, we conducted scalability experiments up to 256 documents (Appendix B.1, Table 8). The findings are twofold: a model trained on 64 documents fails to extrapolate to 256, with its score collapsing from 39.75 to 0.10 EM. However, by continuing to finetune this checkpoint on 128 documents, the model becomes robust at these wider contexts, achieving a strong 37.25 EM on 256 documents. This result proves that souping is not limited by an architectural bottleneck but can be effectively adapted to handle vast contexts. This capability is particularly relevant for applications like retrieval-augmented generation (RAG), where a retriever provides a variable but bounded number of documents (e.g., 20-100) for synthesis. Our results show that document souping provides a practical and highly scalable paradigm for this task. By aligning the finetuning strategy with the expected operational parameters of the RAG system, our method enables robust reasoning over large, dynamic document sets.
4.4 Souping Amplifies Caching Benefits for Efficient Inference
Document souping dramatically improves inference efficiency by enabling a modular caching strategy. Our latency benchmarks, presented in Table 6, demonstrate the substantial speedups achieved by this approach, which excludes the one-time, offline cost of encoding documents. While gains are modest for short sequences, the benefit grows significantly with context length. For a 32k token document, for example, loading the cached state and processing the query takes only 65.8 ms, compared to the 963.2 ms required for the standard concatenation method. This speedup is amplified by the modularity of the souping approach. Because each document’s state is cached independently, any arbitrary subset of documents can be composed on-the-fly at inference time without re-computation. This provides a crucial advantage over monolithic caching schemes, where changing even a single document in a retrieved set would require reprocessing the entire context. This flexibility is particularly powerful for retrieval-augmented generation (RAG) systems, where different queries retrieve unique combinations of documents, making our method a highly efficient solution for dynamic, large-scale reasoning.
@@ 4.4 Souping Amplifies Caching Benefits for Efficient Inference
[…]
@@ 5 Related Work
Transformers (Vaswani et al., 2017) remain the dominant architecture for modeling long-range dependencies, but their $\mathcal{O}(L^{2})$ attention cost and $\mathcal{O}(L)$ memory growth in the KV cache hinder scalability. Many works explore architectural modifications to overcome this, including sparse attention (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020), linearized attention (Katharopoulos et al., 2020), and chunked processing with memory compression (Dai et al., 2019; Wu et al., 2022; Xiao et al., 2024). Retrieval-augmented generation methods, such as RAG (Lewis et al., 2020) offloads long-context memory to external sources.
Structured State Space Models (SSMs) offer an alternative path with fundamentally different scaling characteristics. Models like S4 (Gu et al., 2021, 2022) introduced linear-time computation via structured recurrences. Mamba (Gu and Dao, 2024) builds on this with input-dependent gating, while Mamba2 (Dao and Gu, 2024) unifies SSMs and attention through structured state space duality, yielding substantial speedups on long sequences with strong downstream performance. Recent work further demonstrates the utility of Mamba-based models for dense and retrieval-augmented tasks, including dense passage ranking (Zhang et al., 2024) and long-document retrieval in RAG (Cao et al., 2025).
Model and Representation Souping.
Model souping refers to merging parameters across finetuned checkpoints to improve robustness and generalization without retraining (Wortsman et al., 2022; Tang et al., 2024). More recently, several works have explored souping at the level of internal representations rather than weights. State Soup (pióro2024statesoupincontextskill) stores in-context skills as recurrent states and shows that they can be retrieved and linearly mixed for task transfer. PICASO (Liu et al., 2025) derives a permutation-invariant operator for composing SSM states as an approximation to full concatenation. We extend these works by focusing on souping hidden states corresponding to disjoint document chunks, and study when such aggregated representations can support factual, multi-document reasoning on real downstream benchmarks. Our work characterizes both the regimes in which document-level state souping is effective and the limitations of souping strategies.
@@
Efficient ingestion of long contexts has been addressed through sparse or structured attention mechanisms (Child et al., 2019; Zaheer et al., 2020), memory compression across chunks (Dai et al., 2019; Wu et al., 2022), and retrieval-based pipelines (Lewis et al., 2020). For SSMs, constant-memory recurrence and linear computation make them naturally suited to chunk-wise streaming. However, to our knowledge, no prior work has explicitly studied post-hoc merging of hidden states across chunks for joint reasoning.
@@
Recurrent models have long supported state-passing across time steps, and early works explored implicit mixing of information through recurrent transitions (Grossberg, 2013). Souping differs in that it investigates explicit aggregation of intermediate hidden states produced in parallel, rather than temporal chaining of recurrent updates.
@@
Recent works explored compressing text by adaptive merging KV caches (Wang et al., 2024), and cache eviction policies (Zhang et al., 2023). While related in spirit, these approaches focus on attention-based caches, whereas our work targets SSM hidden states.
@@
In the vision domain, token merging has been studied as a means of compressing long sequences. TCFormer (Zeng et al., 2022) clusters tokens across space for progressive downsampling. Other work merges tokens dynamically based on content (Bolya et al., 2022; Kim et al., 2024), or selects value/key pairs to retain or evict during inference (Wan et al., 2024). These ideas echo our motivation to aggregate compact, composable representations, though in our case applied to document states within state space models.
Table 6: Inference latency with cached states. Souping amplifies caching by allowing any subset of documents to be dynamically composed without re-computation. The initial encoding cost is excluded.
Doc SeqLen Concat Doc + Ques (ms) Cache States + Ques (ms) Speedup
[…]
@@ 6 Conclusion
We study document souping, a method for merging hidden states across independently encoded documents in structured state space models (SSMs), and analyze its behavior on downstream tasks. By aggregating layer-wise representations through simple commutative operations, souping enables modular, parallel ingestion of long-context inputs while supporting accurate downstream inference. Our experiments demonstrate that Mamba2 models trained with souping are capable of multi-hop reasoning, sparse retrieval, and generalization across document and segment scales, often matching or outperforming traditional concat-based finetuning. Overall, our findings highlight state souping as a lightweight and effective strategy for long-context reasoning in SSMs, and we hope it provides a foundation for broader adoption and further exploration.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Data Formatting
We use $\mathbin{\oplus}$ to denote string concatenation, $\langle\mathrm{eos}\rangle$ to mark end-of-sequence, and $\langle\mathrm{DOC\_SEP}\rangle$ to separate documents. Algorithm 1 summarizes the procedure for constructing $x$ under both modes.
Algorithm 1 Input Formatting for Concat- vs. Souping-data
0: Documents $\{d_{1},\ldots,d_{k}\}$, question $q$, answer $a$, flag $\mathrm{soup}\in\{\mathrm{true},\mathrm{false}\}$
0: Formatted input $x$
@@
2:for $i=1$ to $k$ do
3:  if $\mathrm{soup}$ then
4:   $x\leftarrow x\mathbin{\oplus}d_{i}\mathbin{\oplus}\langle\mathrm{DOC\_SEP}\rangle$
5:  else
[…]
7:  end if
8:end for
[…]
@@ A.3 Pseudo-code
0: Document set $\{d_{1},\dots,d_{k}\}$, query $q$
0: Predicted answer $\widehat{y}$
1: Initialize empty list of layerwise states $\{H^{(l)}\}_{l=1}^{L}$
2:for each document $d_{i}$ do
3:  Compute hidden states: $\{h_{i}^{(1)},\dots,h_{i}^{(L)}\}\leftarrow\mathrm{SSM}_{\theta}(d_{i})$
4:  for $l=1$ to $L$ do
5:   Optionally normalize: $\tilde{h}_{i}^{(l)}\leftarrow h_{i}^{(l)}/\|h_{i}^{(l)}\|$
6:   Append $\tilde{h}_{i}^{(l)}$ to $H^{(l)}$
7:  end for
8:end for
9:for $l=1$ to $L$ do
10:  Pool document states: $h_{\text{soup}}^{(l)}\leftarrow\mathrm{pool}(H^{(l)})$
11:  Optionally normalize: $h_{soup}^{(l)}\leftarrow h_{soup}^{(l)}/\|h_{soup}^{(l)}\|$
12:end for
13: Predict: $\widehat{y}\leftarrow\mathrm{SSM}_{\theta}\left(q\mid\{h_{\text{soup}}^{(l)}\}_{l=1}^{L}\right)$
A.4 Implementation and Training Details
This section provides a comprehensive overview of the hyperparameters, software, and hardware used in our experiments.
A.4.1 Training Hyperparameters and Compute Resources
To ensure reproducibility, all experiments were conducted with the random seed fixed to 42 and PyTorch’s deterministic mode enabled (deterministic=True). For distributed training, we employ DeepSpeed ZeRO-2, which enables efficient extraction of the complete hidden states required for our souping mechanism. Our optimization strategy is based on the setup from prior work on Mamba. The specific hyperparameters are as follows:
  • Optimizer: AdamW
  • Learning Rates: The 2.7B Mamba2 model is trained with a learning rate of $5\times 10^{-5}$, while the 8B Mamba2 model uses $1\times 10^{-5}$.
  • AdamW Betas: We use $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
  • Learning Rate Schedule: A cosine decay schedule was used with a linear warm-up for the first 10% of training steps.
  • Gradient Clipping: Gradients were clipped at a maximum norm of 1.0.
  • Gradient Accumulation: We used 120 accumulation steps for most configurations. For the Mamba2-2.7B model on HotpotQA, this was set to 128 steps.
Experiments were conducted on clusters equipped with NVIDIA H100 and H200 GPUs.
A.5 Decoding Hyperparameters
We optimized decoding strategies for both Concat and Soup /w Average using a subset of HotpotQA comprising 30K training examples and 3K validation examples, each containing 5 documents (2 gold, 3 distractors). A grid search was performed over temperature $\{0.3,\ 0.5,\ 0.7,\ 0.9\}$, top-$p$ $\{0.5,\ 0.7,\ 0.9,\ 0.95\}$, and top-$k$ $\{5,\ 10,\ 20,\ 30,\ 40\}$.
Soup achieved best validation performance with temperature =0.3=0.3, top-$p=0.5$, and top-$k=20$, while Concat performed optimally with temperature =0.5=0.5, top-$p=0.95$, and top-$k=30$. All subsequent experiments adopted these respective configurations for each method.
Table 7: HotpotQA results for Mamba-2 2.7B with 22 gold documents and n2n-2 distractors. Cells show EM / F1. Underline marks models tested on the same docs they are trained on.
Model Evaluation: $2$ gold + (n2)(n{-}2) distractors